OpenStudy (anonymous):

1/4x-5=x a)-20/3 b) -15/3 c) 20/3 d) 15/3

OpenStudy (anonymous):

1/4x-5=x | - 1/4x 0.25x-5=x | - 0.25x -5 = 0.75x | :(-5) 1 = -0,15x | :(-0.15) 6.67 = x a) -20/3 = 6.67